# Robust Scene Change Detection Using Visual Foundation Models and Cross-Attention Mechanisms

### v2

@@ I Introduction
Given the limitations of current methods, a more robust approach to scene change detection is needed. Our method addresses these gaps by leveraging a visual foundational model as the backbone network for its robust feature extraction capabilities and integrating cross-attention to register the features. This combination allows for accurate handling of correspondences and mis-correspondences between image pairs, effectively mitigating the impact of photometric and viewpoint changes, and leading to better generalization, as shown in Fig. 1.
@@ II Related Works
One of the most significant challenges in SCD, especially for robotics, is handling viewpoint variations. Researchers often insert feature comparators between encoders and decoders to register features across different aligned images. For instance, SSCDNet [21] employs correlation layers to address viewpoint differences by establishing feature registration, while C3PO [27] proposes multiple subtraction branches to classify changes by learning each type of change separately. Alternatively, some methods treat change detection and feature registration as independent tasks, using optical flow labels to assist in change detection [11, 16].
@@
Compared to methods that classify or recognise the semantic meaning of changes, our approach is orthogonal in its focus on robustly detecting changes under significant viewpoint variations. We leverage a visual foundational model for robust feature extraction and introduce a full-image cross-attention mechanism to effectively handle viewpoint differences between image pairs. By freezing the backbone network during training, we retain the generality of dense foundational features, enhancing the reliability of change detection. Our method is complementary to classification-based approaches and can be integrated with them to address both geometric and semantic challenges in scene change detection.
@@ IV-A Image Encoding:
The DINOv2 [14] backbone is designed to produce all-purpose visual features and is constructed based on the Vision Transformer (ViT) model [5]. Because of its ability to the visual representation, we do not train or fine-tune it to our datasets. Instead, we freeze the smallest one (21M parameters) and use it to develop this model.
The frozen backbone will generate rich features in every image patch (14$\times$14 pixels). So an image $\in\mathbb{R}^{H\times W\times 3}$ will extract to a dense feature $F\in\mathbb{R}^{h\times w\times f}$, where $H$ equals $14\times h$ and $W$ equals $14\times w$.
@@ IV-C Change Mask Prediction:
First, a 3$\times$3 convolution layer is formed to half the number of features from the advanced feature matrix extracted by the image comparator. Second, for segmentation prediction, a 1x1 convolution layer is formed to decode the dense signals into two channels: change and unchanged. Lastly, an upsampling layer is applied to upsample the predicted mask $\in\mathbb{R}^{h\times w\times 2}$ to the target prediction $\in\mathbb{R}^{H\times W\times 2}$.
@@ V-A Datasets
Many change detection datasets are publicly available for benchmarking. CDnet2012 [6] and CDnet2014 [28] released a series of videos to detect changes in outdoor CCTV cameras. ChangeSim [15] recorded drone videos in simulated warehouses to find out artificial changes. These datasets provide enormous images but are limited to a few scenes. VL-CMU-CD [1] is a dataset aiming to update large-scale autonomous vehicle navigation maps and provide many more city scenes, making it a challenging change detection dataset. PSCD [21] gave hundreds of panoramic image pairs for semantic change detection tasks at different places. Both VL-CMU-CD and PSCD have fewer image pairs compared to CDnet2012, CDnet2014 and ChangeSim, but they have more scene to evaluate performance of change detection methods. Thus, we choose VL-CMU-CD and PSCD to evaluate our methods. The number of image pairs and scenes of all datasets are listed in Tab. I.
@@ V-B Evaluation Metric:
Following previous methods [1, 2, 27], we use the F1-score, the harmonic mean of precision and recall, as the evaluation metric. For each image pair in the VL-CMU-CD and PSCD, we compute an F1-score for a predicted change mask. Then, we average the scores from the test sets.
@@ V-C Implementation Details
We followed the training setting from C3PO [27], using the Adam optimizer [9], 0.0001 initial learning rate, and the cosine learning-rate decay strategy. We use the weighted softmax cross-entropy loss function during training, and the weights for the change and unchanged classes are 0.975 and 0.025, respectively. The significant difference between the change and unchanged weight is because most change objects in the datasets only take a small fraction of a whole image. About the training hardware, we use one NVIDIA A100 Tensor Core GPU to train with batch size 4.
@@ VI-A Viewpoint Robustness
We compare a series of baselines in Tab. III and report corresponding performance on the aligned and unaligned VL-CMU-CD and aligned PSCD datasets. For the TransCD [29] and C-3PO [27], we evaluate results by their providing pre-trained weight. We trained DR-TANet [2] and CDNet [22] on our VL-CMU-CD training set, as they do not provide pre-trained weights. Meanwhile, we report CSCDNet [21] result from the paper of C3PO [27] for reference.
@@ VII Conclusion
By effectively managing correspondences between image pairs, our method outperformed existing approaches and proved particularly effective in scenarios involving geometric changes. This robust solution is applicable in autonomous driving, urban planning, environmental monitoring, and surveillance. Future work will focus on further model enhancements and incorporating additional contextual information to improve detection accuracy.

### v3

@@ I Introduction
Given the limitations of current methods, we propose a more robust approach to change detection. Our method addresses these gaps by leveraging a visual foundational model as the backbone network for its robust feature extraction capabilities and integrating cross-attention to register the features. This combination allows for accurate handling of correspondences and mis-correspondences between image pairs, effectively mitigating the impact of photometric and viewpoint changes, and leading to better generalization, as shown in Fig. 1.
@@ II Related Works
A major challenge in SCD, especially for robotics, is handling viewpoint variations. Researchers often insert feature comparators between encoders and decoders to register features across different aligned images. For instance, SSCDNet [21] employs correlation layers to address viewpoint differences by establishing feature registration, while C3PO [27] proposes multiple subtraction branches to classify changes by learning each type of change separately. Alternatively, some methods treat change detection and feature registration as independent tasks, using optical flow labels to assist in change detection [11, 16].
@@
Compared to methods that classify or recognize the semantic meaning of changes, our approach is orthogonal in its focus on robustly detecting changes under significant viewpoint variations. We leverage a visual foundational model for robust feature extraction and introduce a full-image cross-attention mechanism to effectively handle viewpoint differences between image pairs. By freezing the backbone network during training, we retain the generality of dense foundational features, enhancing the reliability of change detection. Our method is complementary to classification-based approaches and can be integrated with them to address both geometric and semantic challenges in scene change detection.
@@ IV-A Image Encoding:
The DINOv2 [14] backbone is designed to produce all-purpose visual features and is constructed based on the Vision Transformer (ViT) model [5]. Because of its strong visual representation capabilities, we do not train or fine-tune it to our datasets. Instead, we freeze the smallest one (21M parameters) and use it to develop this model.
The frozen backbone will generate rich features in every image patch (14$\times$14 pixels). Thus, an image $\in\mathbb{R}^{H\times W\times 3}$ will be transformed into a dense feature $F\in\mathbb{R}^{h\times w\times f}$, where $H$ equals $14\times h$ and $W$ equals $14\times w$.
@@ IV-C Change Mask Prediction:
First, a 3$\times$3 convolution layer is formed to halve the number of features from the advanced feature matrix extracted by the image comparator. Second, for segmentation prediction, a 1x1 convolution layer is formed to decode the dense signals into two channels: change and unchanged. Lastly, an upsampling layer is applied to upsample the predicted mask $\in\mathbb{R}^{h\times w\times 2}$ to the target prediction $\in\mathbb{R}^{H\times W\times 2}$.
@@ V-A Datasets
Many change detection datasets are publicly available for benchmarking. CDnet2012 [6] and CDnet2014 [28] released a series of videos to detect changes in outdoor CCTV cameras. ChangeSim [15] recorded drone videos in simulated warehouses to identify artificial changes. These datasets provide enormous images but are limited to a few scenes. VL-CMU-CD [1] is a dataset aiming to update large-scale autonomous vehicle navigation maps and provide many more city scenes, making it a challenging change detection dataset. PSCD [21] provides hundreds of panoramic image pairs for semantic change detection tasks in different locations. Both VL-CMU-CD and PSCD have fewer image pairs compared to CDnet2012, CDnet2014 and ChangeSim, but they contain more diverse scenes for evaluating change detection methods. Thus, we choose VL-CMU-CD and PSCD to evaluate our methods. The numbers of image pairs and scenes from all datasets are listed in Tab. I.
@@ V-B Evaluation Metric:
Following previous methods [1, 2, 27], we use the F1-score, the harmonic mean of precision and recall, as the evaluation metric. For each image pair in the VL-CMU-CD and PSCD, we compute the F1-score for a predicted change mask. Then, we average the scores from the test sets.
@@ V-C Implementation Details
We followed the training setting from C3PO [27], using the Adam optimizer [9], 0.0001 initial learning rate, and the cosine learning-rate decay strategy. We use the weighted softmax cross-entropy loss function during training, and the weights for the change and unchanged classes are 0.975 and 0.025, respectively. The significant difference between the change and unchanged weight is because most change objects in the datasets only take a small fraction of a whole image. Regarding the training hardware, we used one NVIDIA A100 Tensor Core GPU to train with batch size 4.
@@ VI-A Viewpoint Robustness
We compare a series of baselines in Tab. III and report their respective performance on the aligned and unaligned VL-CMU-CD and PSCD datasets. For the TransCD [29] and C-3PO [27], we evaluate results by their providing pre-trained weight. We trained DR-TANet [2] and CDNet [22] on our VL-CMU-CD training set, as they do not provide pre-trained weights. Meanwhile, we report CSCDNet [21] result from the paper of C3PO [27] for reference.
@@ VII Conclusion
By effectively managing correspondences between image pairs, our method outperformed existing approaches and demonstrated strong performance in scenarios involving geometric changes. This robust solution is applicable in autonomous driving, urban planning, environmental monitoring, and surveillance. Future work will focus on further model enhancements and the incorporation of additional contextual information to improve detection accuracy.